OpenStudy (anonymous):

solve for X = 2= square root (2x−5)−square root(x−2) 

OpenStudy (anonymous):

\[Solve for X: 2=\sqrt{2x-5}-\sqrt{x-2}\]

mathslover (mathslover):

\[\huge{2=\sqrt{2x-5}-\sqrt{x-2}}\] \[\huge{2^2=(\sqrt{2x-5}-\sqrt{x-2})^2}\] \[\huge{4=2x-5+x-2-2\sqrt{2x-5}\sqrt{x-2}}\] \[\huge{4=3x-7-2\sqrt{(2x-5)(x-2)}}\] \[\huge{11=3x-2\sqrt{2x^2-4x-5x+10}}\] \[\huge{11=3x-2\sqrt{2x^2-9x+10}}\]

mathslover (mathslover):

can u do further ?

OpenStudy (anonymous):

how did you get step #3?

OpenStudy (anonymous):

wouldnt it be 2x-5-x-2? thank you for the answer btw. I'm just trying to understand the steps :)

OpenStudy (anonymous):

What he did at step 3 is basically do binomial multiplication.

OpenStudy (anonymous):

oh i see. how would i then solve for x after the last step?

OpenStudy (anonymous):

\[\huge(-\frac{11 - 3x}{2})^{2} = (\sqrt{2x^{2} - 9x + 10})^{2}\]

OpenStudy (anonymous):

Can you finish this?

OpenStudy (anonymous):

I'm sorry i don't quite understand haha. I'm completely stumped

OpenStudy (anonymous):

\[\huge\frac{9x^{2} - 66x + 121}{4} = 2x^{2} - 9x + 10\]@mathslover Did I do it right?

OpenStudy (anonymous):

@mathslover look ^ did I do it right?

mathslover (mathslover):

u did right

OpenStudy (anonymous):

Alright. Now from what I did above: \[9x^{2} - 66x + 121 = 4(2x^{2} - 9x + 10)\]\[9x^{2} - 66x + 121 = 8x^{2} - 36x + 40\]\[x^{2}- 30x + 81\]Can you solve form here?

OpenStudy (anonymous):

That last line should equal 0.

OpenStudy (anonymous):

got it (x-27)(x-3)=0. so 27 and 3

OpenStudy (anonymous):

thanks again calcmathlete

OpenStudy (anonymous):

Yeah. Now plug it back in for x and see if both solutions work.

OpenStudy (anonymous):

and @mathslover

OpenStudy (anonymous):

only 27 works

OpenStudy (anonymous):

Wait, before you finish, be sure to plug 27 and 3 back in to check for extraneous solutions.

OpenStudy (anonymous):

Yeah. That means x = 27 and ≠ 3

OpenStudy (anonymous):

Got it. thanks again!